I need an equation to find how many quarts in 16 pints?
Found 2 solutions by stanbon, lynnlo:
Answer by stanbon(75887)

1 quart = 2 pints
-----
1 pint = (1/2)quart
-------------------------
I need an equation to find how many quarts in 16 pints?
16*1 pint = 16(1/2)quarts
16 pint = 8 quarts
======================
Cheers,
Stan H.
=========================

Answer by lynnlo(4176)

DIVIDE 2PINTS BY 16 WHICH WILL EQUAL 8
AN EQUATION:2 DIVIDED BY 16=8